OpenStudy (anonymous):

A natural number \(y\) is defined as the sum of the digit of the number \( x \), and \( z \) as the sum of the digits of the number \( y \). How many natural numbers \( x \) satisfy the equation \( x+y+z = 60\)?

OpenStudy (anonymous):

wth!?!?! i am so not getting the solution

OpenStudy (anonymous):

lol :D you asked for a hard one ;)

OpenStudy (anonymous):

lol i am not counting it this time LOL

OpenStudy (anonymous):

x must be less than 60, right?

OpenStudy (anonymous):

so we can count! :-D

OpenStudy (anonymous):

There you go IShaan :D

OpenStudy (anonymous):

intuition mwahahaha

OpenStudy (anonymous):

where'd it go?

OpenStudy (lollylau):

What if the question was x + y + z = 1000...

OpenStudy (anonymous):

You don't need to be so happy :P

OpenStudy (anonymous):

i won't even dare to solve it then

OpenStudy (lollylau):

is it called the digital root or sth like that?

OpenStudy (anonymous):

program a program and let the computer do it

OpenStudy (anonymous):

yes exactly, as my teacher would say this CAT/GRE level problems are generally not meant to be generalize but only particularize.

OpenStudy (lollylau):

*these aren't digital roots - search definition in google :)

OpenStudy (lollylau):

anyone still counting?

OpenStudy (lollylau):

wait i think i have an answer

OpenStudy (anonymous):

I was answering to Ishaan.

OpenStudy (lollylau):

lets say there is some x y z that satifisfies x+y+z = 60.

OpenStudy (mr.math):

Please no hints!

OpenStudy (lollylau):

adding 0s doesn't affect the answer.

OpenStudy (anonymous):

i think it's ok to count LOL may be there are not many numbers.

OpenStudy (anonymous):

44, 47

OpenStudy (lollylau):

so if we add 0s in x (or write), that doesn't really affect the answer

OpenStudy (lollylau):

how about 440, 470, 404, 407, 704, 70010000002001?

OpenStudy (lollylau):

have i done it wrong(ly)?

OpenStudy (lollylau):

that's a way to get around the problem :D

OpenStudy (mr.math):

Let \(x=10a+b\), where \(a\le 5\) , \(b\le 4\). Case1: \(a+b\le 9\), then \[x+y+z=10a+b+a+b+a+b=60 \implies 4a+b=20\] The integers solutions are (5,0), (4,4), (3,8).

OpenStudy (anonymous):

and yeah 50 too

OpenStudy (mr.math):

Our first three solutions are \(x=50, 44\) and \(38\).

OpenStudy (anonymous):

oh 38 as well

OpenStudy (anonymous):

Only 1 left.

OpenStudy (anonymous):

50,44,47,38

OpenStudy (lollylau):

it turns out that im wrong :D:D:D:D:D:D:D:D

OpenStudy (anonymous):

Magnum opus Ishaan & Mr.Math :)

OpenStudy (lollylau):

what makes you think of this question???

OpenStudy (anonymous):

hmm lollylau i think it's more about your approach towards the problem, you need a simplistic approach for gre i guess you can't make it complicated and let it eat up your time

OpenStudy (anonymous):

what should be the last number?? hmmmmmmmmm

OpenStudy (lollylau):

program: enter range x to y enter sum s save x as x save y as y A if x>y, stop add up all the numbers of x(value of n*10^x,...n.) save it as b add up again save as c add a b c if a+b+c=s, jot it down (in some way) x=x+1 goto A if x

OpenStudy (lollylau):

the last bit 'if x' shouldn't be there.

OpenStudy (lollylau):

i meant add up x b c

OpenStudy (lollylau):

...

OpenStudy (anonymous):

31 numbers ,0 、1、2、3、..........30right?

OpenStudy (mr.math):

38 is not correct, I made a mistake up there.

OpenStudy (anonymous):

My apologies, there is actually 3 solutions of this form.